OpenStudy (anonymous):

two angles are supplementary. the measure of one angel is 58 degrees greater than the measure of the other angle. find the measures of the angles. So i just got put into honors math. can someone explain to me what this even means?

OpenStudy (anonymous):

supplementry means 180 let the two angles x and y x = 58+y 58+y + y = 180 2y = 180-58 2y =122 y = 122/2 = 61

OpenStudy (anonymous):

x = 58 + 61 = 119

OpenStudy (anonymous):

so...the angles are 61 and 119

OpenStudy (anonymous):

thank you. i am wondering how you got 119. please explain.

OpenStudy (anonymous):

x = 58+y

OpenStudy (anonymous):

we have the value of y

OpenStudy (anonymous):

oh thank you!